OpenStudy (anonymous):

Can anyone help? We need to solver for y. I'm not sure how to do it, so i'd like help :) 12^(y-2)=20

OpenStudy (johnweldon1993):

*Hint...logarithm \[\huge \ln12^{y- 2} = (y - 2)\ln12\] So work with that now... \[\large (y - 2)\ln12 = 20\] Can you solve from there?

OpenStudy (johnweldon1993):

sorry....ln20 on the other side too! \[\large (y - 2)\ln12 = \ln20\]

OpenStudy (anonymous):

Next you would divide both sides by ln12? So it would become y-2=ln(20)/ln(12)?

OpenStudy (johnweldon1993):

Correct....and then the last step would be add 2 to both sides... \[\large y = \frac{ln20}{ln12} + 2\]

OpenStudy (anonymous):

We had to round in to the nearest tenth, so we end up with Y=3.2? is that correct?

OpenStudy (johnweldon1993):

That's right!

OpenStudy (anonymous):

Oh my god thank you so much!

OpenStudy (johnweldon1993):

No problem man!